Use the Pythagorean theorem to determine the length of the unknown side of the triangle,(b) determine the perimeter of the triangle(c) determine the area of the triangle.
Side A=12 in
side B=15 in

Answer by oberobic(2304)

Assuming side C is the hypotenuse.
C^2 = A^2 + B^2
C^2 = 144 + 225 = 369
C = sqrt(369) = 19.2093727
.
Therefore, the perimeter = 12+15+19.2093727.
.
Done.